# Multi-agent Embodied AI: Advances and Future Directions

### v1

@@ 3.3 Generative Model Based Methods
Despite their progress, traditional learning methods often suffer from limited representational capacity, making them inadequate for handling complex scenarios such as multi-goal or multi-task decision making. These methods typically struggle with poor learning performance, inefficient use of external prior knowledge, and low sample efficiency. To address these challenges, researchers have begun exploring the use of generative models, which offer stronger representational capabilities, to enable more efficient and flexible embodied intelligence [53, 197]. The most straightforward application of generative models is to serve as an end-to-end controller by directly outputting executable actions, thereby transferring their internal priors and pretrained knowledge into embodied systems. However, many embodied intelligence tasks involve fine-grained, low-level behaviors, which generative models often struggle to understand and control effectively. Inspired by the principles of hierarchical learning, a growing body of work has adopted generative models for generating high-level plans, while relying on traditional control or learning-based methods to execute low-level skills. Beyond direct decision-making, generative models are also widely used to support the learning process. Applications include integrating sensory information for perception, generating reward functions, and synthesizing data to improve sample efficiency.
@@
To enhance the utilization of prior knowledge and the processing of complex, multimodal observations during decision-making, a growing body of research has explored using generative models, particularly large VLMs, as direct decision-makers, as shown in Figure 13(a). Early approaches typically involved formatting inputs such as states and goals into natural language and feeding them into pretrained generative models, enabling action inference without additional training [29, 83]. However, due to the limited scope of scenarios covered during pretraining, these models often struggle to generalize to the wide range of real-world decision-making tasks. Consequently, more studies have focused on fine-tuning generative models with domain-specific data [216, 77, 38, 86, 122, 79], or integrating them with additional networks [36, 155, 12] to enable more effective fusion of domain knowledge and generative capabilities.
@@
Due to the limited capability of generative models in understanding low-level controls for many tasks, and the substantial resources and costs required for domain-specific fine-tuning, directly employing pre-trained generative models for high-level task planning presents a simpler and more widely applicable approach. Leveraging the reasoning [201] and reflection capabilities [213, 103, 74, 76] inherent in pre-trained generative models, these models can effectively generate goal-oriented plans through iterative reasoning and reflection processes upon receiving inputs such as environmental descriptions, states, and objectives [78, 162, 182], as shown in Figure 13(b). For example, given a high-level task like “pour a glass of water and place it on the table”, an LLM can decompose the task into a clear sequence of actionable steps: (S1) pick up the cup from the table, (S2) locate the water dispenser, (S3) align the cup with the dispenser’s spout, (S4) activate the dispenser and wait until the cup is filled, and (S5) place the filled cup back on the table. Meanwhile, safety in embodied AI planning has garnered increasing attention in recent years [244, 221].
@@
Furthermore, since interactions in real physical environments are costly and errors in robotic manipulation can lead to physical damage or system failures, employing generative models for data generation to enhance sample efficiency is another crucial approach in assisting policy learning. Learning world models based on generative model architectures like Transformer and Diffusion Models, and subsequently generating data via these world models, represents a mainstream method to boost sample efficiency [73, 82, 246, 247, 121, 193], as shown in Figure 13(e). In higher-level language-abstracted environments, some methods directly use pre-trained generative models as world models for future trajectory predictions [131]. Another class of approaches, rather than learning explicit world models, utilizes generative models to create diverse simulation environments, generating additional data through interactions between agents and simulated environments [198, 191, 219, 236].
@@ 3.4 Benchmarks
Embodied-Agent-Interface (EAI) [99] 2024
@@
  • ALFRED [158] is a benchmark designed to evaluate embodied agents’ ability to ground and execute free-form natural language instructions through sequences of physical interactions. Built on the AI2-THOR simulator, it includes 120 visually and functionally diverse household scenes, where agents are tasked with completing goals involving navigation, object manipulation, and irreversible state changes. These tasks span seven categories of household activities and involve 58 distinct object types. To support learning and evaluation, the dataset provides 25,743 human-authored instructions paired with 8,055 expert demonstrations. Agents perceive the environment from egocentric RGB-D inputs and act using a predefined set of 13 discrete low-level actions. Interaction targets are defined through pixel-wise masks, which guide object-level behaviors. Task performance is evaluated primarily based on goal completion, and the benchmark includes all code, annotations, and environment assets necessary to support reproducible research.
  • RoboTHOR [35] is a sim-to-real platform for embodied AI that provides paired simulated and physical environments designed for consistent cross-domain evaluation. It includes 75 training and validation scenes, along with 24 held-out test scenes (14 for test-dev and 10 for test-standard), all constructed from a modular asset library that supports flexible reconfiguration and expansion. Agents interact with the environment via a unified AI2-THOR API and can be remotely deployed on LoCoBot robots, whose noisy dynamics are closely matched to those in simulation. The benchmark is centered on semantic navigation tasks, with performance evaluated using success rate, SPL, and path efficiency in both virtual and physical settings. To promote reproducible research, all code, assets, and remote-access infrastructure are publicly released.
@@
  • EAI [99] is a standardized benchmark for evaluating LLM-driven decision-making in embodied environments. It formalizes task goals using linear temporal logic and structures the decision-making process into four core modules: goal interpretation, sub-goal decomposition, action sequencing, and transition modeling. Each module is evaluated using fine-grained metrics such as grammar correctness, hallucination frequency, step completeness, and ordering accuracy. The benchmark is implemented on top of BEHAVIOR and VirtualHome, and incorporates automated sub-task checkers as well as PDDL-based planners to enable systematic evaluation and reproducible comparison across 18 LLMs in translating natural language instructions into executable plans.
@@
  • PhysBench [32] is a large-scale benchmark designed to evaluate VLMs on their understanding of physical concepts in real-world scenarios. It consists of 10,002 multiple-choice instances that interleave video, image, and text modalities, covering four core domains: physical object properties, object relationships, scene understanding, and physical dynamics. These domains are further divided into 19 fine-grained sub-task types, enabling detailed assessment of model capability across diverse aspects of physical reasoning. The dataset is constructed from a combination of real-world video recordings, web-sourced media, and high-fidelity physical simulations. To support fair and reproducible evaluation, PhysBench includes an online platform with withheld test answers, allowing standardized comparison across 75 representative VLMs.
@@ 4.1 Multi-agent Control and Planning
In MAS, control-based methods remain a fundamental approach for achieving high-precision, real-time decision-making under task constraints. Early methods modeled MAS as a single agent and performed centralized control and planning [159, 6]. However, these approaches face significant scalability challenges. To address this, some methods adopt distributed strategies that control each agent in the MAS independently [178], making them more suitable for large-scale agent systems. Nevertheless, such decentralized methods often struggle with resolving conflicts between agents. To overcome these limitations, EMAPF [72] proposed a grouped multi-agent control framework. It dynamically clusters agents based on their spatial proximity, applying centralized control within each group while ensuring that inter-group control remains independent, as shown in Figure 15. This enables efficient coordination among large teams of aerial robots.
@@ 4.2 Learning for Multi-agent Interaction
Control-based methods in MAS, much like in single-agent scenarios, still face challenges such as high computational overhead and limited generalization. As a result, learning-based approaches remain crucial in multi-agent embodied AI. However, unlike single-agent settings, learning-based methods in MAS must additionally address unique challenges, including asynchronous decision-making, heterogeneous agents, and open multi-agent environments.
@@
In multi-agent embodied intelligence, challenges such as communication delays and hardware heterogeneity across agents often disrupt synchronized interactions and feedback from the real environment, making effective policy learning under asynchronous decision-making a significant challenge. To address this, ACE [224] introduces the concept of macro-actions, where a macro-action serves as a centralized goal for the entire MAS. Individual agents then make multiple asynchronous decisions based on this goal, with delayed feedback from the environment provided only after the macro-action is completed. To facilitate policy learning in this setup, ACE employs a macro-action-based MAPPO algorithm, as shown in Figure 16. This approach has proven effective and has inspired several other works tackling asynchronous decision-making challenges [210, 211].
@@ 4.3 Generative Models based Multi-agent Interaction
Although algorithms designed for asynchronous decision-making, heterogeneous agents, and open environments have advanced learning-based methods in multi-agent embodied AI, challenges persist due to the diversity of cooperative behaviors and the partial or missing nature of observational data, leading to low exploration efficiency and increased learning complexity. To address these issues, generative models have emerged as a powerful tool for enhancing decision-making in multi-agent embodied systems. They can introduce prior knowledge to facilitate explicit task allocation and leverage their strong information-processing capabilities to enable inter-agent communication and observation completion for distributed decision-making. Furthermore, MAS may not only involve autonomous agents but also humans. By harnessing the language understanding and generation capabilities of generative models, human-agent interaction and collaboration can be significantly improved, marking a unique and critical application in this domain. Lastly, the significantly larger exploration space in multi-agent settings makes the challenge of sample efficiency even more pronounced than in single-agent scenarios. Consequently, data augmentation techniques based on generative models are essential for improving data efficiency in multi-agent embodied AI.
@@
To address the challenges posed by the diverse collaborative behaviors in multi-agent embodied AI, an increasing number of studies have begun leveraging the prior knowledge and reasoning capabilities of pre-trained generative models to explicitly assign distinct tasks to different agents, significantly reducing the individual exploration space of each agent. For example, SMART-LLM [85] utilizes a pre-trained language model to first decompose a given task into multiple parallel subtasks and then groups agents based on their capabilities. Tasks are subsequently assigned to each group accordingly, enabling, for instance, the division of cleaning areas among different robotic vacuum cleaners, as illustrated in Figure 19. This paradigm of task decomposition and assignment has become a mainstream planning strategy in multi-agent embodied AI [161, 30, 187]. Building on this, a series of works integrates task assignment and execution. These approaches first use generative models for task allocation, then continue using the same or similar models to execute the tasks, adjusting the allocation based on execution outcomes to form a more complete workflow [183, 107, 207, 28]. However, the aforementioned methods generally focus on task assignment based solely on generative reasoning, often overlooking the dependencies among subtasks. For instance, in the task “retrieve the wrench from a closed box,” the subtask “open the box” must be completed before “take the wrench out of the box.” To better capture such dependencies, recent works have begun exploring the use of subtask dependency graphs to enhance task allocation [199, 132]. In addition to explicit task decomposition, another line of work employs centralized generative models to produce global decisions. These are then implicitly distributed across agents by having each agent imitate the behavior generated by the centralized model, thus achieving implicit task assignment [105].
@@
Using generative models for multi-agent embodied task decomposition and planning leverages the prior knowledge and reasoning capabilities acquired during pretraining to meet the demands of real-world collaborative tasks. However, centralized task planning and allocation can compromise the flexibility and scalability of cooperation, often requiring frequent recalls for adjustment. Therefore, exploring the coordination among multiple generative models is essential. Unlike independent methods in MARL that often fail in complex tasks, the outstanding perception and reasoning capabilities of generative models make it feasible to deploy multiple LLM-based agents that independently perform decision-making and policy evaluation effectively [248, 194, 14, 64, 172, 189, 228]. For example, [64] equips each agent with its own pre-trained generative model. These models leverage their strong information processing capabilities to communicate with other agents, either to complete missing observations or to request assistance, thus facilitating decentralized collaboration, as illustrated in Figure 20. However, challenges such as credit assignment and policy conflicts still hinder the effectiveness of fully distributed collaborative architectures. Therefore, similar to practices in MARL, some studies have begun to incorporate an additional centralized generative model to evaluate the decisions made by distributed generative models, enhancing their overall decision-making capabilities [248, 64]. Furthermore, by introducing a shared global LLM-based reflector to assess the contribution of each individual during collaboration, effective credit assignment in multi-LLM cooperation can also be achieved. [14]. With appropriately designed system topologies and hierarchical collaboration frameworks, generative model-based distributed decision making is capable of scaling to large-scale systems including up to thousands of agents [240, 142].
@@
Efficient collaboration between humans and AI agents (robots) has long been a critical objective within AI research [2]. Human-AI coordination [185], closely related to research fields such as human-AI interaction (HAI) [180] and human-robot interaction (HRI) [134], focuses on enhancing teamwork between humans and artificial agents to effectively achieve complex tasks. Cooperative MARL, recognized for its robust problem-solving capabilities, provides promising avenues for improving human-AI collaboration across diverse user groups. However, traditional RL approaches often fall short in fully capturing the intricacies and variability inherent in human behaviors. Recently, with the rise of multi-modal large-scale models, particularly LLMs, which inherently involve human-in-the-loop training processes, researchers have begun leveraging the extensive knowledge embedded in these models to design sophisticated and adaptive strategies for human-AI collaboration [242, 168, 64, 46, 62, 104, 5, 106, 54] as illustrated in Figure 21.
In multi-LLM-based agent scenarios, collaboration between humans and AI agents can be partially facilitated by substituting human participants for communication partners or team leaders [242, 168, 64]. Nevertheless, this approach often underutilizes the strengths humans bring to the collaboration process. Given humans’ superior capacity for understanding, reasoning, and adapting, recent methods have emphasized active human involvement to address limitations inherent in LLM-based agents [46]. For example, during navigation tasks, agents encountering uncertainty can proactively seek missing perceptual or decisional information from humans through natural language queries, allowing humans to offer guidance without directly engaging in detailed task planning or execution [51]. Additionally, LLM-based agents can proactively infer human intentions based on verbal and behavioral cues, facilitating more intuitive, flexible, and effective human-AI collaboration without explicit communication requests [188]. Furthermore, these agents can autonomously adapt and refine their collaborative behaviors over time, evolving through interactions without the need for explicit human instructions, thereby supporting sustained and effective long-term human-AI collaboration [104].
@@
Due to the high sample efficiency of model-based approaches, applying world models to multi-agent collaborative learning has always been a significant topic. However, modeling the interactions between agents and inferring the global state from local observations remains a key difficulty in leveraging world models to improve multi-agent collaboration efficiency. Early work attempted to model the dynamics of two collaborative robots simultaneously, achieving some success [90], as shown in Figure 22(a). However, as the number of collaborative agents increases, this approach quickly becomes inefficient. Through the use of generative models such as VAE and Transformer for local observation fusion, decoupling of global and local modeling, and auto-regressive trajectory prediction, it is now possible to model the dynamic collaboration of MAS using world models [157, 184, 7, 110], as shown in Figure 22(b). One approach addresses these challenges in the context of urban traffic by constructing a generative world model that simulates large-scale heterogeneous agents under physical and social constraints, using destination- and personality-conditioned dynamics to promote realistic and safe multi-agent behaviors [249]. Complementarily, another line of work focuses on physically embodied agents operating with egocentric RGBD observations, introducing a diffusion-based reconstruction process to establish a shared representation of the environment and incorporating intention inference to facilitate implicit coordination without direct communication [243], as shown in Figure 22(c).

### v2

@@ 3.3 Generative Model Based Methods
Despite their progress, traditional learning methods often suffer from limited representational capacity, making them inadequate for handling complex scenarios such as multi-goal or multi-task decision making. These methods typically struggle with poor learning performance, inefficient use of external prior knowledge and low sample efficiency. To address these challenges, researchers have begun exploring the use of generative models, which offer stronger representational capabilities, to enable more efficient and flexible embodied intelligence [54, 204]. The most straightforward application of generative models is to serve as end-to-end controllers by directly outputting executable actions, thereby transferring their internal prior knowledge into embodied agents. However, many embodied tasks involve fine-grained low-level behaviors which generative models often struggle to understand and control effectively. Inspired by the principles of hierarchical learning, a growing body of work has adopted generative models for generating high-level plans, while relying on traditional control or learning-based methods to execute low-level skills. Beyond direct decision-making, generative models are also widely used to support the learning process. Applications include integrating sensory information for perception, generating reward functions, and synthesizing data to improve sample efficiency.
@@
To enhance the utilization of prior knowledge and the processing of complex, multi-modal observations during decision-making, a growing body of research has explored using generative models, particularly large VLMs, as direct decision-makers, as shown in Figure 13(a). Early approaches typically involve formatting inputs such as states and goals in natural language and feed them into pre-trained generative models, enabling action inference without additional training [30, 84]. However, due to the limited scope of scenarios covered during pre-training, the models often struggle to generalize to the wide range of real-world decision-making tasks. Consequently, more studies have focused on fine-tuning generative models with domain-specific data [223, 78, 39, 88, 125, 80], or integrating them with additional networks [37, 161, 12] to improve the generalizability.
@@
Due to the limited capability of generative models in understanding low-level controls for many tasks, combined with the substantial resources and costs required for domain-specific fine-tuning, directly employing pre-trained generative models for high-level task planning presents a simpler and more widely applicable approach. Leveraging the reasoning [208] and reflection capabilities [220, 106, 75, 77] inherent, pre-trained generative models can effectively make goal-oriented plans through iterative reasoning and reflection processes upon inputs such as environmental descriptions, states and objectives [79, 168, 188], as shown in Figure 13(b). For example, given a high-level task like “pour a glass of water and place it on the table”, an LLM can decompose the task into a clear sequence of actionable steps: (S1) pick up the cup from the table, (S2) locate the water dispenser, (S3) align the cup with the dispenser’s spout, (S4) activate the dispenser and wait until the cup is filled, and (S5) place the filled cup back on the table. Meanwhile, safety in embodied AI planning has garnered increasing attention in recent years [251, 228].
@@
Furthermore, since interactions in real physical environments are costly and errors in robotic manipulation can lead to physical damage or system failures, employing generative models for data generation to enhance sample efficiency is another crucial approach in assisting policy learning [254, 152, 128]. Learning world models based on generative model architectures like Transformer and diffusion Models, and subsequently generating data via these world models, represents a mainstream method to boost sample efficiency [74, 83, 253, 255, 124, 200], as shown in Figure 13(e). In higher-level language-abstracted environments, some methods directly use pre-trained generative models as world models for future trajectory predictions [135]. Another class of approaches, rather than learning explicit world models, utilizes generative models to create diverse simulation environments, generating additional data through interactions between agents and simulated environments [205, 198, 226, 243].
@@ 3.4 Benchmarks
Embodied-Agent-Interface [101] 2024
@@
  • ALFRED [164] is a benchmark for evaluating embodied agents’ ability to ground and execute free-form natural language instructions through physical interactions. Built on the AI2-THOR simulator, it includes 120 visually and functionally diverse household scenes, where agents are tasked with completing goals involving navigation, object manipulation, and irreversible state changes. These tasks span seven categories of household activities and involve 58 distinct object types. To support learning and evaluation, the dataset provides 25,743 human-authored instructions paired with 8,055 expert demonstrations. Agents perceive the environment from egocentric RGB-D inputs and act using a predefined set of 13 discrete low-level actions. Interaction targets are defined through pixel-wise masks, which guide object-level behaviors. Task performance is evaluated primarily based on goal completion, and the benchmark includes all code, annotations, and environment assets necessary to support reproducible research.
  • RoboTHOR [36] is a sim-to-real embodied benchmark that provides paired simulated and physical environments designed for consistent cross-domain evaluations. It includes 75 training and validation scenes, along with 24 held-out test scenes (14 for test-dev and 10 for test-standard), which all constructed from a modular asset library that supports flexible reconfiguration and expansion. The agents in RoboTHOR interact with the environment via a unified AI2-THOR API, and the policies can be remotely deployed on real LoCoBot robots whose noisy dynamics are closely matched to those in the simulations. RoboTHOR is centered on semantic navigation tasks with performance evaluated using success rate, SPL and path efficiency in both virtual and physical settings. To promote reproducible research, all code, assets and remote-access infrastructure are publicly released.
@@
  • Embodied-Agent-Interface [101] is a standardized benchmark for evaluating LLM-driven decision-making in embodied environments. It formalizes task goals using linear temporal logic and structures the decision-making process into four core modules: goal interpretation, sub-goal decomposition, action sequencing, and transition modeling. Each module is evaluated using fine-grained metrics such as grammar correctness, hallucination frequency, step completeness, and ordering accuracy. The benchmark is implemented on top of BEHAVIOR and VirtualHome, and incorporates automated sub-task checkers as well as PDDL-based planners to enable systematic evaluation and reproducible comparison across 18 LLMs in translating natural language instructions into executable plans.
@@
  • PhysBench [33] is a large benchmark designed for evaluating VLMs on their understanding of physical concepts in real-world scenarios. It consists of 10,002 multiple-choice instances that interleave video, image and text modalities, covering four core domains: physical object properties, object relationships, scene understanding and physical dynamics. These domains are further divided into 19 fine-grained sub-task types, enabling detailed assessment of model capabilities across diverse aspects of physical reasoning. The dataset of PhysBench is constructed from a combination of multiple resources including real-world video recordings, web-sourced media and high-fidelity physical simulations. To support fair and reproducible evaluation, PhysBench includes an online platform with withheld test answers, allowing standardized comparison across 75 representative VLMs.
@@ 4.1 Multi-agent Control and Planning
In MAS, control-based methods remain a fundamental approach for achieving high-precision, real-time decision-making under task constraints. Early methods model MAS as single agents and perform centralized control and planning [165, 6] methods on them. However, these approaches face significant scalability challenges. To address this, some methods adopt distributed strategies that control each agent in the MAS independently [184], making them more suitable for large-scale MAS. Nevertheless, such fully decentralized methods often struggle with resolving conflicts between agents. To overcome these limitations, EMAPF [73] proposes a grouped multi-agent control framework. It dynamically clusters agents based on their spatial proximity, applying centralized control within each group while ensuring that inter-group control remains independent, as shown in Figure 15. This enables efficient coordination among large teams of aerial robots. Alternatively, Swarm-Formation [148] employs a decentralized sequential decision-making scheme: agents are pre-ordered and each agent plans its trajectory based on its predecessor’s plan in turn, thereby eliminating inter-agent conflicts.
@@ 4.2 Learning for Multi-agent Interaction
Control-based methods in MAS, much like in single-agent scenarios, still face challenges such as high computational overhead and limited generalizability. As a result, learning-based approaches remain crucial in constructing multi-agent embodied AI. However, unlike single-agent settings, learning-based methods in MAS must additionally address unique challenges, including asynchronous decision-making, heterogeneous team composition and open multi-agent environments.
@@
In multi-agent embodied systems, challenges such as communication delays and hardware heterogeneity across agents often disrupt synchronized interactions and feedback from the real environment, making effective policy learning under asynchronous decision-making a significant challenge. To address this, ACE [231] introduces the concept of macro-actions, where a macro-action serves as a centralized goal for the entire MAS. Individual agents then make multiple asynchronous decisions based on this goal, with delayed feedback from the environment provided only after the macro-action is completed. To facilitate policy learning in this setup, ACE employs a macro-action-based MAPPO algorithm, as shown in Figure 16. This approach has proven effective and has inspired several other works tackling asynchronous decision-making challenges [217, 218].
@@ 4.3 Generative Models based Multi-agent Interaction
Although algorithms designed for asynchronous decision-making, heterogeneous agents and open environments have advanced learning-based methods in embodied MAS, challenges persist due to the diversity of cooperative behaviors and the partial or missing nature of observations, leading to low exploration efficiency and increased learning complexity. To address these issues, generative models have emerged as a powerful tool for enhancing decision-making in embodied MAS. They can introduce prior knowledge to facilitate explicit task allocation and leverage their strong information-processing capabilities to enable inter-agent communication and observation completion for distributed decision-making. Furthermore, MAS may not only involve autonomous agents but also humans. By harnessing the language understanding and generation capabilities of generative models, human-agent interaction and collaboration can be significantly improved, marking a unique and critical application in this domain. Lastly, the significantly larger exploration space in multi-agent settings makes the challenge of sample efficiency even more pronounced than in single-agent scenarios. Consequently, data augmentation techniques based on generative models are essential for improving data efficiency in multi-agent embodied AI.
@@
To address the challenges posed by the diverse collaborative behaviors in embodied MAS, an increasing number of studies have begun leveraging the prior knowledge and reasoning capabilities of pre-trained generative models to explicitly assign distinct tasks to different agents, significantly reducing the individual exploration spaces of each agent. For example, SMART-LLM [87] utilizes a pre-trained language model to first decompose a given task into multiple parallel sub-tasks and then group agents based on their capabilities. The sub-tasks are subsequently assigned to each group accordingly, enabling multi-agent decentralized task executions. For instance, different robotic vacuum cleaners are each allocated a distinct cleaning area, as illustrated in Figure 19. This paradigm of task decomposition and assignment has become a mainstream planning strategy in embodied MAS [167, 31, 193]. Building on this, a series of works integrates task assignment and execution. These approaches first use generative models for task allocation, then continue using the same or similar models to execute the tasks, adjusting the allocation based on execution outcomes to form a more complete workflow [189, 110, 214, 29]. However, the aforementioned methods generally focus on task assignment based solely on generative reasoning, often overlooking the dependencies among sub-tasks. For instance, in the task “retrieve the wrench from a closed box”, the sub-task “open the box” must be completed before “take the wrench out of the box”. To better capture such dependencies, recent works have been focused on exploring the use of sub-task dependency graphs to enhance task allocation [206, 136]. In addition to explicit task decomposition, another line of work employs centralized generative models to produce global decisions. These are then implicitly distributed across agents by having each agent imitate the behavior generated by the centralized model, thus achieving implicit task assignment [108].
@@
Using generative models for multi-agent embodied task decomposition and planning leverages the prior knowledge and reasoning capabilities acquired during pretraining to meet the demands of real-world collaborative tasks. However, centralized task planning and allocation can compromise the flexibility and scalability of cooperation, often requiring frequent recalls for adjustment. Therefore, exploring the coordination among multiple generative models is essential. Unlike independent methods in MARL that often fail in complex tasks, the outstanding perception and reasoning capabilities of generative models make it feasible to deploy multiple LLM-based agents that independently perform decision-making and policy evaluation effectively [256, 201, 14, 65, 178, 196, 235]. For example, [65] equips each agent with a unique pre-trained generative model. These models leverage their strong information processing capabilities to communicate with other agents, either to complete missing observations or to request assistance, thus facilitating decentralized collaboration, as illustrated in Figure 20. However, challenges such as credit assignment and policy conflicts still hinder the effectiveness of fully distributed collaborative architectures. Therefore, similar to practices in MARL, some studies have begun to incorporate an additional centralized generative model to evaluate the decisions made by distributed generative models, enhancing their overall decision-making capabilities [256, 65]. Furthermore, by introducing a shared global LLM-based reflector to assess the contribution of each individual during collaboration, effective credit assignment in multi-LLM cooperation can also be achieved. [14]. With appropriately designed system topologies and hierarchical collaboration frameworks, generative model-based distributed decision making is capable of scaling to large-scale systems including up to thousands of agents [247, 146].
@@
Efficient collaboration between humans and embodied agents has long been a critical objective of research [2]. Human-AI coordination [191], closely related to research fields such as human-AI interaction (HAI) [186] and human-robot interaction (HRI) [138], focuses on enhancing teamwork between humans and robots to effectively achieve complex tasks. Cooperative MARL, recognized for its robust problem-solving capabilities, provides promising avenues for improving human-AI collaboration across diverse user groups. However, traditional RL approaches often fall short in fully capturing the intricacies and variability inherent in human behaviors. Recently, with the rise of multi-modal large models such as LLMs which inherently involve human-in-the-loop training processes, researchers have begun leveraging the extensive knowledge embedded in them to design sophisticated and adaptive strategies for human-AI collaboration [249, 174, 65, 47, 63, 107, 5, 109, 55] as illustrated in Figure 21.
In scenarios including multiple LLM-based agents, collaboration between humans and agents can be partially facilitated by substituting human participants for communication partners or team leaders [249, 174, 65]. Nevertheless, this approach often under-utilizes the strengths that humans bring to the collaboration process. Given humans’ superior capacities for understanding, reasoning, and adapting, recent methods have emphasized active human involvement to address limitations inherent in LLM-based agents [47]. For example, during navigation tasks, agents encountering uncertainty can proactively seek missing perceptual or decisional information from humans through natural language queries, allowing humans to offer guidance without directly engaging in detailed task planning or execution [52]. Additionally, LLM-based agents can proactively infer human intentions based on verbal and behavioral cues, facilitating more intuitive, flexible, and effective human-AI collaboration without explicit communication requests [194]. Furthermore, these agents can autonomously adapt and refine their collaborative behaviors over time, evolving through interactions without the need for explicit human instructions, thereby supporting sustained and effective long-term human-AI collaboration [107].
@@
Due to the high sample efficiency of model-based RL approaches, applying world models to enhance multi-agent collaborative learning has always been a significant topic. However, modeling the interactions between agents and inferring the global state from local observations remains a key difficulty in leveraging world models to improve multi-agent collaboration efficiency. Early work attempted to model the dynamics of two collaborative robots simultaneously, achieving some success [92], as shown in Figure 22(a). However, as the number of collaborative agents increases, this approach quickly becomes inefficient. Through the use of generative models such as VAE and Transformers for local observation fusion, decoupling of global and local modeling and auto-regressive trajectory prediction, it is now possible to model the dynamic collaboration of MAS using world models [163, 190, 7, 113], as shown in Figure 22(b). One approach addresses these challenges in the context of urban traffic by constructing a generative world model that simulates large-scale heterogeneous agents under physical and social constraints, using destination- and personality-conditioned dynamics to promote realistic and safe multi-agent behaviors [257]. Complementarily, another line of work focuses on physically embodied agents operating with egocentric RGBD observations, introducing a diffusion-based reconstruction process to establish a shared representation of the environment and incorporating intention inference to facilitate implicit coordination without direct communication [250], as shown in Figure 22(c).